A runner starts from rest and accelerates at a uniform rate 10m/s in the time interval of 5s, running on a circular track of radius 50m. Determine the tangential acceleration.
Answer #1

Tangential acceleration

at = (v/t)

at = 10/ 5 = 2 m/s^2

=========

Comment in case any doubt please rate my answer....